Grade 9Mechanics

A a force of 25 Newton acts on a body at rest for 0.2 seconds and the force of 70 Newton acts for the next 0.1s second in opposite direction if final velocity of body is 5 metre per second the mass of the body is

surendra surendra
7 Years agoGrade 9

1 Answer

Arun
7 Years ago
Let the mass of the body be m.
Case:1
Force of 25N acts on the body for 0.2s
We know that F=ma
F=ma
F=m[(v-u)/t]
25=m[(v-0)/0.2]
5=mv
v=5/m
This velocity acts as initial velocity for the next case.
Case:2
Force of 70N acts for the next 0.1s in Opposite direction.
Initial velocity=5/m
Final velocity=-5 {negative as because in opposite direction}
Force=-70N
F=ma
-70=m[(-5-5/m)/0.1]
-7=m[-5-5/m]
-7=-5m-5
-2=-5m
m=2/5=0.4Kg

Hence mass of the body is 0.4Kg.